# MammAlps: A multi-view video behavior monitoring dataset of wild mammals in the Swiss Alps

### v1

@@ 2 Related Works
Table 1: Prominent and publicly available video datasets of wild animals behavior monitoring. *Multimodal data is available but it is not used for an action recognition benchmark. MammAlps is available at 10.5281/zenodo.15040901.
@@ 3.4 Multi-view long-term event understanding Benchmark: B2
Modalities Spe.($\uparrow$) ActY.($\uparrow$) ActN.($\uparrow$) Avg.($\uparrow$)
V 0.487 0.428 0.458 0.453
S 0.414 0.188 0.171 0.211
A 0.223 0.207 0.161 0.184
V+S 0.457 0.399 0.375 0.394
A+S 0.334 0.262 0.257 0.270
V+A 0.503 0.475 0.463 0.472
V+A+S 0.482 0.452 0.417 0.438
Table 4: Hierarchical action recognition from multimodal data (B1). mAP for joint task prediction from multimodal data using VideoMAE as the base model [47]. V: video clips; A: audio spectrograms; S: segmentation map clips; ActY.: Activities; ActN.: Actions; Spe.: Species; Avg.: overall per-class average. Note: the ‘V’ row, corresponds to the last row of Tab. 3.
Training task $r$ Cont. Len. Spe.($\uparrow$) ActY.($\uparrow$) Met. Cond.($\uparrow$) Indiv.($\uparrow$) Avg. ($\uparrow$)
@@
All 14 4096 0.343 0.483 0.653 0.478 0.476
All 11 8192 0.439 0.450 0.634 0.593 0.498
Table 5: mAP for long-term event understanding from the multi-view events (B2). All models use the transformer encoder from ViT-Base. "$r$": ToME [7] reduction factor. A larger reduction factor leads to more patches being merged at the frame level and fewer video tokens; "Cont. Len.": context length: number of tokens per sample; ActY.: Activities; Spe.: Species.; Met. Cond.: Meteorological Conditions; Indiv.: Number of individuals categories.; Avg.: overall per-class average.
For both benchmarks B1 and B2, we report the mean average precision (mAP) averaged over the label categories of each task, which is a convenient metric to compare tasks that are either multi-label or multi-class. When applicable, for joint predictions on all tasks, we report the mAP averaged over all label categories of all tasks in column ‘Avg.’. Models for benchmarks B1 and B2 were trained with four and eight A100 GPUs, respectively.
@@ 4.1 B1: Multi-modal species and behavior recognition
When considering only the video modality, VideoMAE leads to improved results for all tasks when considering the joint task prediction (Tab. 3). Multi-modal results indicate that combining the audio and video modalities improves the performance over the video-only model ($+0.019$ mAP), with an overall class-average mAP of $0.472$ (Tab. 4). However, in our baseline model, the reference segmentation map clips did not improve over their video-only or video-audio counterparts, but they did increase the performance of the audio-only model ($+0.059$ mAP) suggesting that this modality contains distinct information relevant to the tasks. More details, baselines and results per class can be found in the Supplementary Materials.
@@ 4.2 B2: Multi-view long-term event understanding
The best joint recognition performance (average per-class mAP of $0.498$) was achieved with a ToME [7] reduction factor ($r$) at the frame level of 11, yielding between $65$ and $390$ tokens per video depending on their duration (Tab. 5). When $r=11$, the overall mAP is slightly higher ($+0.022$) than when $r=14$ (yielding between $29$ and $174$ tokens per video) but not on all tasks.
@@ 9.1 Multimodal VideoMAE Implementation details
Figure 7: Multimodal Video Transformer implementation for B1. Note that the transformer backbone is similar to both B1 and B2. In B2, the backbone is followed by four classification heads instead of the three depicted here, one for each of the classification tasks.
9.2 Baseline performances.
@@ 9.3 Models performance per class
We report model performances (F1-scores and average precisions) per class (Tab. 9, Tab. 10, Tab. 11 and 12). The advantage of reporting the mAP (or AP when considering single classes) is that the metric better represents the area under the curve as it computes the precision over multiple thresholds, and it can be equally applied to multi-class and multi-label problems. To compute the F1-score, we used a threshold of 0.5 on the softmax and sigmoid outputs for multi-class and multi-label tasks, respectively.
@@ 10.5 Models performance per class
We report F1-scores and average precisions per class computed similarly as for B1 (Tab. 15 and 14). We show the results when using a ToME [7] reduction factor of $r=14$ and $r=11$, and all types of positional encodings ($Cam_{ID}$, $\Delta T_{event}$, $Source$).

### v2

@@ 2 Related Works
Table 1: Prominent and publicly available video datasets of wild animals behavior monitoring. *Multimodal data is available but it is not used for an action recognition benchmark. MammAlps is available at 10.5281/zenodo.15040900.
@@ 3.4 Multi-view long-term event understanding Benchmark: B2
Mod. Spe.($\uparrow$) ActY.($\uparrow$) ActN.($\uparrow$) Avg.$\pm$±std ($\uparrow$)
V 0.495 0.436 0.452 0.453 $\pm$ 0.002
S 0.441 0.234 0.172 0.230 $\pm$ 0.014
A 0.223 0.212 0.172 0.192 $\pm$ 0.004
V+S 0.466 0.409 0.384 0.403 $\pm$ 0.006
A+S 0.385 0.312 0.276 0.303 $\pm$ 0.012
V+A 0.473 0.484 0.466 0.473 $\pm$ 0.017
V+A+S 0.531 0.485 0.437 0.466 $\pm$ 0.007
Table 4: Hierarchical action recognition from multimodal data (B1). mAP for joint task prediction from multimodal data using VideoMAE as the base model [47] averaged across three runs. Mod.: Modalities; V: video clips; A: audio spectrograms; S: segmentation map clips; ActY.: Activities; ActN.: Actions; Spe.: Species; Avg.: overall per-class average.
Training task $r$ Cont. Len. Spe.($\uparrow$) ActY.($\uparrow$) Met. Cond.($\uparrow$) Indiv.($\uparrow$) Avg.$\pm$±std ($\uparrow$)
@@
All 14 4096 0.415 0.479 0.618 0.499 0.489 $\pm$ 0.033
All 11 8192 0.446 0.481 0.594 0.543 0.500 $\pm$ 0.004
Table 5: mAP for long-term event understanding from the multi-view events (B2). Results are averaged across three runs for the joint task predictions. All models use the transformer encoder from ViT-Base. "$r$": ToME [7] reduction factor. A larger reduction factor leads to more patches being merged at the frame level and fewer video tokens; "Cont. Len.": context length: number of tokens per sample; ActY.: Activities; Spe.: Species.; Met. Cond.: Meteorological Conditions; Indiv.: Number of individuals categories.; Avg.: overall per-class average.
For both benchmarks B1 and B2, we report the mean average precision (mAP) averaged over the label categories of each task, which is a convenient metric to compare tasks that are either multi-label or multi-class. When applicable, for joint predictions on all tasks, we report the mAP averaged over all label categories of all tasks in column ‘Avg.’. For the joint task predictions, we report the mAP averaged across three runs using different seeds (along with the std. for the ‘Avg.‘ column). Models for benchmarks B1 and B2 were trained with four and eight A100 GPUs, respectively.
@@ 4.1 B1: Multi-modal species and behavior recognition
When considering only the video modality, VideoMAE leads to improved results for all tasks when considering the joint task prediction (Tab. 3). Multi-modal results indicate that combining the audio and video modalities improves the performance over the video-only model ($+0.020$ mAP), with an overall class-average mAP of $0.473$ (Tab. 4). However, in our baseline model, the reference segmentation map clips did not improve over their video-only or video-audio counterparts, but they did increase the performance of the audio-only model ($+0.111$ mAP) suggesting that this modality contains distinct information relevant to the tasks. More details, baselines and results per class can be found in the Supplementary Materials.
@@ 4.2 B2: Multi-view long-term event understanding
The best joint recognition performance (average per-class mAP of $0.500$) was achieved with a ToME [7] reduction factor ($r$) at the frame level of 11, yielding between $65$ and $390$ tokens per video depending on their duration (Tab. 5). When $r=11$, the overall mAP is slightly higher ($+0.011$) than when $r=14$ (yielding between $29$ and $174$ tokens per video) but not on all tasks.
@@ 6 Conclusion and outlook
Changes from CVPR published version and arXiv v1: After submission, we noticed that a few audio files were not correctly aligned with the corresponding video. We fixed the issue, which had little to no impact on performance (Tab. 4). The released data on Zenodo already reflects this correction: 10.5281/zenodo.15040900. During this process, we also ran more replicates. Consequently, results for Tab. 4 and for the joint task predictions of Tab. 5 have been updated to show results for three runs. We now report the mAP averaged across three runs along with the standard deviation for the average mAP to provide a better estimate of the performance variation across runs. We report the standard deviation for all tasks in the Supplementary Material. Results in the text have been updated accordingly.
@@ 9.1 Multimodal VideoMAE Implementation details
Figure 7: Multimodal Video Transformer implementation for B1. The transformer backbone is similar to both B1 and B2. In B2, the backbone is followed by four classification heads instead of the three depicted here, one for each of the classification tasks.
9.2 Baseline performance and variability.
@@
Mod. Spe. ActY. ActN. Avg.
V 0.495 $\pm$ 0.020 0.436 $\pm$ 0.016 0.452 $\pm$ 0.014 0.453 $\pm$ 0.002
S 0.441 $\pm$ 0.050 0.234 $\pm$ 0.027 0.172 $\pm$ 0.010 0.230 $\pm$ 0.014
A 0.223 $\pm$ 0.014 0.212 $\pm$ 0.010 0.172 $\pm$ 0.002 0.192 $\pm$ 0.004
V+S 0.466 $\pm$ 0.005 0.409 $\pm$ 0.009 0.384 $\pm$ 0.018 0.403 $\pm$ 0.006
A+S 0.385 $\pm$ 0.012 0.312 $\pm$ 0.014 0.276 $\pm$ 0.017 0.303 $\pm$ 0.012
V+A 0.473 $\pm$ 0.013 0.484 $\pm$ 0.036 0.466 $\pm$ 0.011 0.473 $\pm$ 0.017
V+A+S 0.531 $\pm$ 0.018 0.485 $\pm$ 0.014 0.437 $\pm$ 0.011 0.466 $\pm$ 0.007
Table 9: Variability of the mAPs for the joint task predictions of B1 Mean and standard deviation are computed after training the model three times with different seeds.
@@ 9.3 Models performance per class
We report model performances (F1-scores and average precisions) per class (Tab. 10, Tab. 11, Tab. 12 and 13). The advantage of reporting the mAP (or AP when considering single classes) is that the metric better represents the area under the curve as it computes the precision over multiple thresholds, and it can be equally applied to multi-class and multi-label problems. To compute the F1-score, we used a threshold of 0.5 on the softmax and sigmoid outputs for multi-class and multi-label tasks, respectively.
@@ 10.4 Camera-views ablation
$r$ Cont. Len. Spe. ActY. Met. Cond. Indiv. Avg.
14 4096 0.415 $\pm$ 0.084 0.479 $\pm$ 0.032 0.618 $\pm$ 0.034 0.499 $\pm$ 0.019 0.489 $\pm$ 0.033
11 8192 0.446 $\pm$ 0.059 0.481 $\pm$ 0.032 0.594 $\pm$ 0.035 0.543 $\pm$ 0.033 0.500 $\pm$ 0.004
Table 15: Variability of the mAPs for the joint task predictions of B2 Mean and standard deviation are computed after training the model three times with different seeds. "$r$": ToME [7] reduction factor. A larger reduction factor leads to more patches being merged at the frame level and fewer video tokens; "Cont. Len.": context length: number of tokens per sample; ActY.: Activities; Spe.: Species.; Met. Cond.: Meteorological Conditions; Indiv.: Number of individuals categories.; Avg.: overall per-class average.
@@ 10.5 Models performance per class
We report F1-scores and average precisions per class computed similarly as for B1 (Tab. 17 and 16). We show the results when using a ToME [7] reduction factor of $r=14$ and $r=11$, and all types of positional encodings ($Cam_{ID}$, $\Delta T_{event}$, $Source$).